Sales are higher in the warmer portion of the year, so April-Sept sales are usually double Oct-Mar sales. How is this calculated, if only have a total amount of sales from Jan-Aug?
Answer #1

In order to be able to solve the above question, we will have to use the following assumptions :

1. Month-wise sales in the winter portion are equal and month-wise sales in the warmer portion of the year are equal which means that if sale for January is 1000 $, then sale for February will also be 1000 $ and sales for any month of the warmer portion of the year will be 2000 $.

2. We will have to assume sales from Jan-Aug to be any number. I have assumed it to be 26,000 $.

Now, let monthly sale for any month in Oct-Mar be x $. Hence sales for any month between Apr-Sept becomes 2x $.

Sales for Jan-Aug = 26,000 $, i.e. (x+x+x+2x+2x+2x+2x+2x) = 26,000 $

Solving the above equation, we get x = 2,000 $.

Hence sales for Oct-Mar = 6x = 2,000 $*6 = 12,000 $

And sales in warmer portion, i.e. Apr to Sept = 2*12,000 $ = 24,000 $

Total sales for the year = 36,000 $.

Note : Students can make alternate assumptions as well.